OpenStudy (anonymous):

What is the equation, in standard form, of a vertical hyperbola with asymptotes at y + 8 = ±2/9(x – 8)?

OpenStudy (agent0smith):

Vertical hyperbola has asymptotes: \[y=\pm \frac{ a }{ b }(x-h)+k\] You have: \[y = ±\frac{ 2 }{ 9 }(x – 8) - 8\] and the standard form of a vert. hyperbola: \[\frac{ (y-k)^2 }{ a^2 }-\frac{ (x-h)^2 }{ b^2 } = 1\]

OpenStudy (anonymous):

what is the y?

OpenStudy (agent0smith):

I'm not sure what you mean...

OpenStudy (agent0smith):

y is a variable like x

OpenStudy (anonymous):

yeah, but in y= ±2/9(x – 8)-8, isthere a value for y?

OpenStudy (anonymous):

to use to convert it to standard form?

OpenStudy (agent0smith):

No, there's no need for y or x values. If you look at my first reply, compare the first two formulas - they give you a, b, h and k.

OpenStudy (agent0smith):

Once you find a, b, h, k by comparing the first two, insert them into the second equation for the standard form.

OpenStudy (anonymous):

k= -8 h=8 a=2

OpenStudy (agent0smith):

Correct, and b?

OpenStudy (anonymous):

9

OpenStudy (agent0smith):

Yep! Now just enter them into: \[\frac{ (y-k)^2 }{ a^2 }-\frac{ (x-h)^2 }{ b^2 } = 1\]

OpenStudy (anonymous):

(y+8)^2/4- (x+8)^2/81=1

OpenStudy (agent0smith):

The (x+8) should be (x-8), since h = 8

OpenStudy (agent0smith):

Change that and it all looks correct.